# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

### v1

@@ 4.3 Instruction-following
LLaMA2 Chat 7B 87.3
LLaMA2 7B + LoRA 85.9
LLaMA2 7B + PiSSA 85.6
LLaMA2 7B + MiLoRA 87.3
Setup
We follow the experiment setup in Wu et al. (2024a) and Wu et al. (2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023). Following previous works, we use Alpaca-Eval v1.0  (Li et al., 2023) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024). We implement our method using open-instruct [^4] https://github.com/allenai/open-instruct.
Hyperparameter
We use the LLM-Adapters hyperparameter setup for all PEFT methods. We truncate samples to a maximum of 768 tokens. During evaluation, we use the last checkpoint with a greedy decoding strategy and a maximum token number of 2048. To avoid the issue of repetition, we set the repetition penalty to 1.1 and set the no repeat n-gram size to 5, following Wu et al. (2024a) and Wu et al. (2024b).
@@ Results
Table 3 presents the win-rate of various PEFT methods. We include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2 Chat 7B as a reference, which have win-rate of 91.2%, 91.8%, and 87.3% respectively. Our method, LLaMA2-7B + MiLoRA outperforms the other two PEFT methods and matches LLaMA2 Chat 7B’s performance, indicating that MiLoRA is a highly effective finetuning method.
@@ 5 Understanding MiLoRA
In this section, we conduct experiments to further understand MiLoRA. For all analysis experiments that do not involve finetuning, we use the LLaMA2-7B model finetuned on MetaMathQA in our main experiments (see Section 4.2). For any experiments that require finetuning except for Section 5.5, we finetune the LLaMA2-7B model on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512. This configuration is chosen to minimize computational costs.
5.1 How do different rank settings influence MiLoRA?
Figure 2: Average score of finetuned LLaMA2-7B model on math reasoning across different ranks.
In Figure 2, we explore the impact of various rank configurations by adjusting $r$ within the set $\{16,32,64,128\}$ and assessing the average accuracy of the finetuned LLaMA2-7B model on math reasoning tasks. Detailed numbers are presented in Table 10 of Appendix B.1. As can be seen, MiLoRA with rank 64 achieves the best performance across all methods and ranks. MiLoRA consistently surpasses PiSSA across all rank configures. However, when compared with LoRA, MiLoRA performs better only at large ranks. This is likely due to the fact that MiLoRA makes fewer changes to the pre-trained parameters compared to LoRA (see detailed discussion in Section 5.3 and 5.4). As a result, when the rank is relatively low, MiLoRA struggles to fully digest the finetuning dataset.
@@ 5.2 Is minor singular component adaptation important?
To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model. Specifically, we initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and evaluate the results on the GSM8K and MATH datasets.
@@
Method Principle Random Minor
GSM8K 60.65 63.23 63.99
MATH 14.60 15.54 16.14
The results, shown in Table 4, demonstrate that initializing with the minor singular components achieves the best performance across both evaluation datasets. This underscores the importance of adapting the minor singular components during the finetuning process. Concurrent with our work, Meng et al. (2024) propose PiSSA, a similar approach to MiLoRA but adapts the principle singular components. We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
@@ 5.3 How does the matrix update $\Delta W$ compare to $W$?
We investigate the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ of different methods, where $\Delta{\bm{W}}$ is the difference between the finetuned and pre-trained weights. Following the analysis method in Hu et al. (2021), we project ${\bm{W}}$ and $\Delta{\bm{W}}$ onto the $r$-dimensional subspace of $\Delta{\bm{W}}$ by computing ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ and ${\bm{U}}^{T}\Delta{\bm{W}}{\bm{V}}$, with ${\bm{U}}$/${\bm{V}}$ being the left/right top $r$ singular-vector matrix of $\Delta{\bm{W}}$. As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix. We use the Frobenius norm as the measurement of magnitude. The results are shown in Table 5. Following Hu et al. (2021), we use the query weight in the middle layer (16th) of the model for analysis. We make two key observations. First, $\Delta{\bm{W}}$ of LoRA and MiLoRA has a similar correlation with ${\bm{W}}$, both slightly stronger compared to a random matrix, indicating that they amplify directions that are not emphasized in ${\bm{W}}$. In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$. This is attributed to PiSSA’s direct optimization of the principal singular components. Second, the amplification factor varies significantly across different methods. Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller. This indicates that MiLoRA has a reduced impact on features already present in ${\bm{W}}$. See Table 11 of Appendix B.2 for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
@@
5.4 Does MiLoRA forget less than baselines?
Table 6: Forgetting loss of various PEFT methods.
@@ 5.4 Does MiLoRA forget less than baselines?
A hypothesis explanation for why MiLoRA outperforms baselines is that pre-trained knowledge is more retained. To test this hypothesis, we follow Kalajdzievski (2024) and use cross-entropy as the metric for measuring forgetting. This is the usual next token prediction loss used when training LLMs, except that the target next token is replaced by the distribution predicted by the pre-trained base model. We evaluate the forgetting metric on LLaMA2-7B model finetuned in Section 4.2 using the WikiText-103 test dataset following (Kalajdzievski, 2024). As shown in Table 6, MiLoRA exhibits the lowest forgetting loss, which is consistent with our hypothesis that MiLoRA makes the least modification to pre-trained knowledge. In contrast, the loss of PiSSA is significantly higher than that of the other methods.
@@ 5.5 Comparison between MiLoRA and PiSSA
Concurrent with our research, Meng et al. (2024) have recently proposed a low-rank adaptation method called PiSSA. At first glance, PiSSA may appear to be quite similar to MiLoRA but adapts the principle singular components. However, there are fundamental differences between PiSSA and MiLoRA that are worth highlighting:
@@
  • Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components. However, our investigation suggests that this claim is likely due to the specific hyperparameters used in their experiments. To verify this, we replicate their experimental setup using the same MetaMathQA 100K dataset and compare the performance of PiSSA, LoRA, and MiLoRA under both PiSSA’s hyperparameters and our own. Specifically, we choose a rank of 128 where PiSSA obtains the best performance in their paper and finetunes for 1 epoch for all setups following PiSSA paper. The results are summarized in Table 7. We have several observations. First, under the PiSSA hyperparameter setup, PiSSA does indeed outperform both LoRA and MiLoRA, consistent with the claims made in the original PiSSA paper. Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective. Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.39 and 4.27 average scores, respectively. In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
We further investigate why the hyperparameter configuration in PiSSA paper is suboptimal. This configuration, as shown Table 9 of Appendix A.3, is from the full finetuning hyperparameters of Alpaca (Taori et al., 2023). As can be seen, it has a combination of small learning rate (2e-5) and large batch size (128). This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023). This problem becomes more severe as PiSSA paper only trains for a single epoch for all experiments.
Figure 4: Validation loss curve of different PEFTs under PiSSA hyperparameter configurations. LoRA and MiLoRA are under-fitting when training with 1 epoch (781 steps).
To validate our assumption, we extract another 120 samples from MetaMathQA to serve as a validation set and then train LoRA, PiSSA, and MiLoRA under the same hyperparameter configuration except for 10 epochs each and recorded their respective validation loss curves. From the results shown in Figure 4, we make two key observations. First, LoRA and MiLoRA are much under-fitting when training with only 1 epoch (781 steps). However, with sufficient training, their validation loss can converge to the same level as that of PiSSA. Second, PiSSA converges faster than the other two methods but quickly overfits. This might be attributed to that PiSSA directly finetunes the essence of weight via the top singular vectors and values. We evaluate the best checkpoint of each methods selected by the validation set, and find that LoRA performs best, while the other two perform similarly (see Table 12 in Appendix B.3), which again confirms that LoRA and MiLoRA are observed worse than PiSSA in the PiSSA paper because they are underfitting.
Hyp. Setup PEFT GSM8K MATH Avg.
PiSSA LoRA 41.47 5.76 23.6
PiSSA 51.25 7.58 29.4
MiLoRA 40.03 5.18 22.6
LLM-Adapters LoRA 56.56 10.84 33.70
PiSSA 51.25 10.38 30.82
MiLoRA 58.61 11.56 35.09
@@ 6 Related Work
After pretraining, large language models are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023). Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of fully finetuning. Previous PEFT methods can be grouped into three lines. The first group is adapter-based, which inserts additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b). However, adapter-based methods introduce additional computation costs during inference. The second category is LoRA-based (Hu et al., 2021; Liu et al., 2024a; Zhang et al., 2023), which models the changes of selected linear weights as low-rank matrices. During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen. The third group is prompt-based (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2022) which adds additional soft prompts to the input tokens. During training, only soft prompts are updated while the pretrained model parameters are fixed. Among these PEFT methods, the LoRA-based approaches are most widely used for LLM finetuning as it does not introduce computation overhead during inference while approximating the performance of fully finetuning.
@@
LoRA (Hu et al., 2021) reparameterizes the weight update with two trainable low-rank matrices, while freezing the pretrained weights. With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks. Theoretically, the linear design of LoRA enables seamless merging of trainable matrices with frozen weights, thereby avoiding extra inference latency. Since then, researchers explore and propose different LoRA variants for PEFT. AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024b), adaptively determine the rank of LoRA module in each weight matrix according to the importance score. The pruned ranks are compensated on the more important positions. The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks. The DoRA (Liu et al., 2024a) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance. VeRA (Kopiczko et al., 2024) reduces the number of trainable parameters compared to LoRA by sharing a pair of low-rank matrices across all layers. Concurrent with our work, PISSA (Meng et al., 2024) proposes to only update the principle singular components. However, their conclusion is empirically problematic due to the sub-optimal hyperparameters.

### v2

@@ 4.1 Experiments on Large Language Model
LLaMA2-Chat-7B 87.3
LLaMA2-7B + LoRA 85.9
LLaMA2-7B + PiSSA 85.6
LLaMA2-7B + MiLoRA 87.3
Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Results are cited from Ivison et al. (2023).
@@ Results
Method VQAv2 GQA VisWiz SQA VQAT POPE MMBench Avg.
Full FT 78.5 61.9 50.0 66.8 58.2 85.9 64.3 66.5
LoRA 79.1 62.9 47.8 68.4 58.2 86.4 66.1 66.9
PiSSA 77.5 60.6 39.2 67.3 54.3 87.0 61.4 63.9
MiLoRA 79.2 62.1 53.3 70.6 58.7 87.9 66.1 68.3
Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks. We directly use checkpoints from Liu et al. (2024a) to reproduce the results of baseline methods.
We report results in Table 1, Table 2 and 3 for commonsense reasoning, math reasoning and instruction following, respectively. For commonsense reasoning, we also include the ChatGPT baseline reported in Liu et al. (2024b) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022). For math reasoning, we add the full finetuning results from Yu et al. (2023) as a reference. For instruction following, we include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2-Chat-7B from Ivison et al. (2023) as a reference.
As can be seen, MiLoRA consistently surpasses all baseline methods across different datasets and LLMs, indicating that MiLoRA is a highly effective finetuning method. Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B, and by 1.1 and 6.5 points on LLaMA3-8B. For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged EM points on LLaMA2-7B, respectively. However, it still underperforms full finetuning by an average of 2.5 EM score. This suggests that the PEFT methods still have room for improvement to fully match the performance of full finetuning. For instruction following, MiLoRA achieves the best results, surpassing LoRA and PiSSA by 1.4 and 2.7 in win-rate scores, respectively.
4.2 Experiments on Vision-Language Model
Models and Datasets
To further examine if MiLoRA can remain competitive on multimodal finetuning tasks, we further conduct visual instruction tuning tasks on LLaVA1.5-7B Liu et al. (2024a), which consists of an LLM, Vicuna-1.5-7B Peng et al. (2023) and a vision encoder, CLIP ViT-L/336px Radford et al. (2021). The visual instruction tuning is performed in a multitask setting, including VQA Hudson and Manning (2019); Marino et al. (2019); Schwenk et al. (2022), OCR Mishra et al. (2019); Sidorov et al. (2020), region-level VQA Krishna et al. (2017); Mao et al. (2016), visual conversation Liu et al. (2024a), and language conversation data. The finetuned models are evaluated on seven different multimodal benchmarks: VQA-v2 Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), SQA Lu et al. (2024), VQAT Singh et al. (2019), POPE Li et al. (2023b), and MMBench Liu et al. (2024c).
Implementation Details
We follow the settings of Liu et al. (2024a) to construct the training data and prompt template. For the hyperparameter configuration, we follow the LoRA configuration provided by Liu et al. (2024a) without tuning for all PEFT methods for a fair comparison. Details can be found in Appendix A.2.
Results
Table 4 shows the experiment results. As can be seen, MiLoRA achieves the best results, outperforming LoRA and PiSSA by 1.4 and 4.4 average accuracy points, respectively. We also note that in this setup, fully fine-tuning is less effective than LoRA. As a result, methods like PiSSA, which aim to closely approximate fully fine-tuning, may lose their advantage. In contrast, MiLoRA improves performance by retaining more of the pre-trained knowledge, thereby circumventing this issue.
@@ 5 Understanding MiLoRA
In this section, we conduct experiments to further understand MiLoRA. By default, we use the finetuned LLaMA2-7B model on math reasoning from our main experiments for analyses.
5.1 Comparison with More PEFTs
To further demonstrate the effectiveness of MiLoRA, we compare MiLoRA with additional popular baselines, such as rsLoRA Kalajdzievski (2023), LoRA+ Hayou et al. (2024), DoRA Liu et al. (2024b), AdaLoRA Zhang et al. (2023), LoRA-GA Wang et al. (2024b). We follow the math and code experiment settings in Wang et al. (2024b). Specifically, we finetune LLaMA2-7B with MiLoRA on MetaMathQA 100K and Code-Feedback 100K for one epoch using rank $8$ and compare MiLoRA with baseline results reported in Wang et al. (2024b). Table 5 shows the results. We report the training cost of all methods relative to LoRA, which requires 3.58 hours of training on the math task using a single NVIDIA-L40 GPU. As can be seen, MiLoRA consistently outperforms all baselines, improving the best PEFT baseline LoRA-GA by 1.1 and 4.6 points on GSM8K and Human-eval, respectively. The training efficiency of MiLoRA is also very high. The time for SVD decomposition in MiLoRA is less than six minutes and negligible when compared with the total training time. Therefore, the training time required of MiLoRA is consistent with that of LoRA, while DoRA and AdaLoRA require additional training time overhead.
Method GSM8K Human-eval Training Cost
LoRA 42.1 14.8 $\times$ 1.00
PiSSA 44.5 16.0 $\times$ 1.00
rsLoRA 45.6 16.0 $\times$ 1.00
LoRA+ 52.1 18.2 $\times$ 1.00
DoRA 53.1 19.8 $\times$ 1.58
AdaLoRA 50.7 17.8 $\times$ 1.07
LoRA-GA 53.6 19.8 $\times$ 1.00
MiLoRA 54.7 24.4 $\times$ 1.00
Table 5: Comparison with more baselines on math and code tasks. Baseline GSM8K and Human-eval results are cited from Wang et al. (2024b).
@@ 5.2 Is Minor Singular Component Adaptation Important?
To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model. Specifically, we finetune LLaMA2-7B on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512 to save computation. We initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and report the evaluation results in Table 6. As can be seen, initializing with the minor singular components achieves the best performance across both datasets. This underscores the importance of adapting the minor singular components during the finetuning process. Concurrent with our work, Meng et al. (2024) propose PiSSA, a similar approach to MiLoRA but adapts the principal singular components. We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
Method principal Random Minor
GSM8K 60.7 63.2 64.0
MATH 14.6 15.5 16.1
@@ 5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
We investigate the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ of different methods, where $\Delta{\bm{W}}$ is the difference between the finetuned and pre-trained weights. Following the analysis method in Hu et al. (2021), we project ${\bm{W}}$ and $\Delta{\bm{W}}$ onto the $r$-dimensional subspace of $\Delta{\bm{W}}$ by computing ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ and ${\bm{U}}^{T}\Delta{\bm{W}}{\bm{V}}$, with ${\bm{U}}$/${\bm{V}}$ being the left/right top $r$ singular-vector matrix of $\Delta{\bm{W}}$. As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix. We use the Frobenius norm as the measurement of magnitude. The results are shown in Table 7. Following Hu et al. (2021), we use the query weight in the middle layer (16th) of the model for analysis. We make two key observations. First, $\Delta{\bm{W}}$ of LoRA and MiLoRA have a similar correlation with ${\bm{W}}$, both slightly stronger than a random matrix, indicating that they amplify directions not emphasized in ${\bm{W}}$. In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$. This is attributed to PiSSA’s direct optimization of the principal singular components. Second, the amplification factor varies significantly across different methods. Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller. This indicates that MiLoRA has a reduced impact on the features which already present in ${\bm{W}}$. See Table 13 of Appendix B for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
@@
5.4 MiLoRA Forgets Less than Baselines
@@
Table 8: Forgetting loss of various PEFT methods.
A hypothesis explanation for why MiLoRA outperforms baselines is that pre-trained knowledge is more retained. To test this hypothesis, we follow Kalajdzievski (2024) and use cross-entropy as the metric for measuring forgetting. This is the usual next token prediction loss used when training LLMs, except that the target next token is replaced by the distribution predicted by the pre-trained base model. We evaluate the forgetting metric on the LLaMA2-7B model finetuned in math reasoning using the WikiText-103 test dataset following Kalajdzievski (2024). As shown in Table 8, MiLoRA exhibits the lowest forgetting loss, which is consistent with our hypothesis that MiLoRA makes the least modification to pre-trained knowledge. In contrast, the loss of PiSSA is significantly higher than that of the other methods.
@@ 5.5 Comparison Between MiLoRA and PiSSA
Hyp. Setup PEFT GSM8K MATH Avg.
PiSSA LoRA 41.5 5.8 23.6
PiSSA 51.3 7.6 29.4
MiLoRA 40.0 5.2 22.6
LLM-Adapters LoRA 56.6 10.8 33.7
PiSSA 51.3 10.4 30.8
MiLoRA 58.6 11.6 35.1
Concurrent with our research, Meng et al. (2024) have recently introduced a low-rank adaptation method called PiSSA. PiSSA shares a similar framework as MiLoRA, but adapts the principal singular components. We argue that PiSSA and MiLoRA are fundamentally different.
@@
Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components. However, our investigation suggests that this claim is likely due to the specific hyperparameters used in their experiments. To verify this, we replicate their experimental setup using the same MetaMathQA 100K dataset and compare the performance of PiSSA, LoRA, and MiLoRA under both PiSSA’s hyperparameters and our own. Specifically, we choose a rank of 128 where PiSSA obtains the best performance in their paper and finetunes for 1 epoch for all setups following PiSSA paper. The results are summarized in Table 9. We have several observations. First, under the PiSSA hyperparameter setup, PiSSA does indeed outperform both LoRA and MiLoRA, consistent with the claims made in the original PiSSA paper. Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective. PiSSA configuration of hyperparameters has a combination of a small learning rate (2e-5) and a large batch size (128). This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023; Biderman et al., 2024). Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.4 and 4.3 average scores, respectively. In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
@@ 6 Related Work
After pretraining, LLMs are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023). Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of full finetuning. Previous PEFT methods can be grouped into three lines: adapter-, LoRA- and prompt-based methods. Adapter-based methods insert additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b). However, they introduce additional computational costs during inference. LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023) model the changes of selected linear weights as low-rank matrices. During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen. Prompt-based methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022) add additional soft prompts to the input tokens. During training, only soft prompts are updated while the pretrained model parameters are fixed. Among these PEFT methods, the LoRA-based approaches are widely used for LLM fine-tuning because they are easy to implement and do not introduce computational overhead during inference.
@@
LoRA (Hu et al., 2021) reparameterizes the weight update with two trainable low-rank matrices, while freezing the pretrained weights. With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks. Since then, researchers have explored and proposed different LoRA variants for PEFT. AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024d), adaptively determine the rank of LoRA module in each weight matrix according to the importance score. The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks. The DoRA (Liu et al., 2024b) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance. Concurrent with our work, PISSA (Meng et al., 2024) and LoRA-GA Wang et al. (2024b) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.